OpenStudy (anonymous):

Write the equation for a line perpendicular to the line y= -2x+5 and containing the point (3,1)

OpenStudy (anonymous):

well the slope of the line given is -2 right? The perpendicular of the line is the opposite reciprocal (opp, and flip it) so the new slope is 1/2 y=1/2x+b our points are (3,1) which are (x,y) we can use these points to plug into the equation above. 1=1/2(3)+b in order to get b... b is -.5 our final equation will be y=1/2x-0.5